Find f+g,f-g,fg, and f/g and their domains.
f(x)=x-3, g(x)=x^2
I am really struggling with this section. Please help.

Answer by stanbon(75887)

Find f+g,f-g,fg, and f/g and their domains.
f(x)=x-3, g(x)=x^2
----
f+g = x^2+x-3
Domain: All Real Numbers
----
f-g = x^2-x+3
All Real Numbers
----
fg = x^2(x-3) = x^3-3x^2
Domain: All Real Numbers
----
f/g = (x-3)/x^2
Domain:: All Real Numbers except x = 0
=============
Cheers,
Stan H.
==============